# Multi-level protein pre-training with Vabs-Net

### v3

@@ 4.4 Molecular Docking
Predicting the binding structure of a small molecule ligand to a protein, a task known as molecular docking, is crucial for drug design (Pei et al., 2023). Equibind (Stärk et al., 2022) and Diffdock (Corso et al., 2023) are the two most commonly used models for this purpose. However, the diffusion-based Diffdock requires substantial computational resources; thus, we employ Equibind (Stärk et al., 2022) to evaluate the efficacy of features from our model. Equibind is trained using features extracted from various structural pre-training models. To be more specific, we trained 4 distinct Equibind models. The first one is Equibind itself. For the other three models, we add features from GearNet (pre-trained by multiview contrast learning), Siamdiff (atom-level), and our Vabs-Net to the node representation of Equibind. Those features are preprocessed before training. It is also important to note that Equibind is a residue-level model. To optimize the utilization of atom-level features from both the Siamdiff model and our own, we implement a single attention layer to aggregate atom features to the residue node representation of Equibind.
Setting For molecular docking, due to time constraints, we adopt a faster and more stable version of Equibind, which can also be seen in the official repository. More detail is in Appendix D. To conduct a fair comparison between our model and other pre-training models, we incorporate ESM features into all the models under consideration. Both Equibind and its variant enhanced with pre-trained features are trained for 300 epochs. We use PDBBind as dataset, which is also used by Equibind (Stärk et al., 2022).
@@ 4.5 ABLATION STUDIES
To analyze the effect of different components, we choose a protein function prediction task(EC) and a binding site prediction task(small molecule(SM)) for ablation study.
We investigate different pre-training strategies and model configurations, with the findings presented in Table 6. (1) A comparison between models No.0 and No.1 demonstrates the significance of atom-level encoding, particularly in binding site prediction tasks where side chain atoms play a crucial role. (2) Upon comparing models No.0 and No.2, it becomes clear that residue-level encoding is also important for enhancing the receptive field and capturing residue-level representations. (3) Furthermore, the comparison between models No.0 and No.3 confirms the effectiveness of our proposed SMPC pre-training method in enhancing residue-level representations. (4) The notable decrease in performance observed after removing the vector edge encoder (No.3 and No.4) underscores the importance of direction encoding, rather than solely encoding distance. (5) While leveraging ESM can enhance performance (No.5 and No.3), the impact is not as significant as that of the vector edge encoder. (6) Increasing the KNN parameter does not yield a substantial improvement in the model’s performance (No.5 and No.3), possibly due to KNN=30 being adequate for capturing most structural motifs (Tateno et al., 1997).
@@
No.
Method SM EC
0
Vabs-Net 64.1 0.914
1
No Atom level 56.3 0.901
2
No Residue level 62.2 0.900
3
No SMPC 63.0 0.898
4
No SMPC,Vector Edge Encoder 60.9 0.879
6
No SMPC,ESM 61.9 0.877
5
No SMPC,knn=90 62.7 0.896
@@ 5 Conclusion
In this study, we introduce the Vabs-Net model with the span mask strategy 3D protein chains pre-training technique, aiming to learn atom-level representation and improve residue-level representation. We have conducted extensive experiments on various task types to assess the effectiveness of our Vabs-Net model and the span mask protein chains pre-training approach. Our Vabs-Net model demonstrates superior performance, surpassing previous state-of-the-art models. In the subsequent phase of our research, we plan to enhance the integration of sequence and structural feature.
6 Impact Statements
@@
Dataset Method
ave.
B277
DT198
ASTEX85
CHEN251
COACH420
scPDB ESM2_150M
18.4
18.8
15.6
17.2
18.5
21.8
GearNet
29.6
28.8
26.2
31.2
29.0
32.8
Siamdiff_atom
26.6
25.2
21.6
28.1
27.8
30.1
@@
Our Dataset ESM2_150M
19.5
19.6
16.6
20.5
18.9
22.0
GearNe
38.9
39.9
35.8
41.0
36.4
41.3
Siamdiff
37.0
37.7
31.0
40.7
35.3
40.3
Vabs-Net 58.0 60.7 51.9 60.0 55.3 63.0
Appendix B More detailed information about model architecture.
@@
Vabs-Net 64.1 0.914
No SMPC 63.0 0.898
No SMPC SO(3) 60.9 0.897
@@
Throughout the fine-tuning process, with the exception of molecular docking, all parameters remained unfrozen, except for ESM features where the ESM features are computed before training. Additionally, in conventional practice, two evolutionary conversation profiles (PSSM and HMM) are traditionally utilized in DNA and RNA binding site prediction (Xia et al., 2021). A more comprehensive description of the model is provided in Table 9. The sizes of DNA and RNA molecules are considerably larger than those of small molecules, necessitating the selection of a k-nearest neighbor (knn) value of 90 to encompass long-range information. In contrast, the pre-training model maintains a knn value of 30. Due to the restricted size of the training set for DNA and RNA binding site prediction, we utilized a model with 256-node dimensions.
@@
Appendix C More detailed information about downstream task experiments.

### v4

@@ 4.4 Molecular Docking
Predicting the binding structure of a small molecule ligand to a protein, a task known as molecular docking, is crucial for drug design (Pei et al., 2023). Equibind (Stärk et al., 2022) and Diffdock (Corso et al., 2023) are the two most commonly used models for this purpose. However, the diffusion-based Diffdock requires substantial computational resources; thus, we employ Equibind (Stärk et al., 2022) to evaluate the efficacy of features from our model. Equibind is trained using features extracted from various structural pre-training models. To be more specific, we trained 4 distinct Equibind models. The first one is Equibind itself. For the other three models, we add features from GearNet (pre-trained by multiview contrast learning), Siamdiff (atom-level), and our Vabs-Net to the node representation of Equibind. Those features are pre-processed before training. It is also important to note that Equibind is a residue-level model. To optimize the utilization of atom-level features from both the Siamdiff model and our own, we implement a single attention layer to aggregate atom features to the residue node representation of Equibind.
Setting For molecular docking, due to time constraints, we adopt a faster and more stable version of Equibind, which can also be seen in the official repository. More detail is in Appendix E. To conduct a fair comparison between our model and other pre-training models, we incorporate ESM features into all the models under consideration. Both Equibind and its variant enhanced with pre-trained features are trained for 300 epochs. We use PDBBind as a dataset, which is also used by Equibind (Stärk et al., 2022).
@@ 4.5 Ablation Studies
To analyze the effect of different components, we choose a protein function prediction task(EC) and a binding site prediction task(small molecule(SM)) for the ablation study.
We investigate different pre-training strategies and model configurations, with the findings presented in Table 6. (1) A comparison between models No.0 and No.1 demonstrates the significance of atom-level encoding, particularly in binding site prediction tasks where side chain atoms play a crucial role. (2) Upon comparing models No.0 and No.2, it becomes clear that residue-level encoding is also important for enhancing the receptive field and capturing residue-level representations. (3) Furthermore, the comparison between models No.0 and No.3 confirms the effectiveness of our proposed SMPC pre-training method in enhancing residue-level representations. (4) The notable decrease in performance observed after removing the vector edge encoder (No.3 and No.4) underscores the importance of direction encoding, rather than solely encoding distance. (5) While leveraging ESM can enhance performance (No.5 and No.3), the impact is not as significant as that of the vector edge encoder. (6) Increasing the KNN parameter does not yield a substantial improvement in the model’s performance (No.6 and No.3), possibly due to KNN=30 being adequate for capturing most structural motifs (Tateno et al., 1997). (7) Our pre-training strategy significantly enhances the model’s performance, as evidenced by the comparisons between No.7 and No.0. (8) To demonstrate the effectiveness of our Sparse Attention Module, we replaced it with the AttMLP module used in PiFold (Gao et al., 2022) in experiment No. 8. The comparison between No.0 and No.8 clearly illustrates the superior performance of our Sparse Attention Module. (9) The results of ablation studies No.9 and No.10 indicate that neither randomly masking protein chains (not necessarily consecutive) nor only masking side chain atoms (while retaining backbone atoms) can completely prevent information leakage.
@@
No. Method SM EC
0 Vabs-Net 64.1 0.900
1 No Atom level 56.3 0.896
2 No Residue level 62.2 0.886
3 No SMPC 63.0 0.876
4 No SMPC,Vector Edge Encoder 60.9 0.874
5 No SMPC,ESM 61.9 0.867
6 No SMPC,knn=90 62.7 0.876
7 No pre-train 61.4 0.826
[…]
9 Span mask side chain atoms 63.6 0.892
10 Randomly mask 63.3 0.884
5 Limitation
Compared to the sequential pre-training dataset (1103M in xTrimoPGLM), our structural pre-training dataset (130K) is much smaller. The AlphaFold database now contains over 100 million structures. Also, our pre-training is limited to single-chain pre-training, neglecting interaction between chains. In our subsequent efforts, we seek to tackle the limitations above.
@@ 6 Conclusion
In this study, we introduce the Vabs-Net model with the span mask strategy 3D protein chains pre-training technique, aiming to learn atom-level representation and improve residue-level representation. We have conducted extensive experiments on various task types to assess the effectiveness of our Vabs-Net model and the span mask protein chains pre-training approach. Our Vabs-Net model demonstrates superior performance, surpassing previous state-of-the-art models. In the subsequent phase of our research, we plan to enhance the integration of sequence and structural features.
Acknowledgments
This study was supported by the National Key Research and Development Program of China (2022YFA1004304).
Impact Statement
@@ Acknowledgments
Dataset Method ave. B277 DT198 ASTEX85 CHEN251 COACH420
scPDB ESM2_150M 18.4 18.8 15.6 17.2 18.5 21.8
GearNet 29.6 28.8 26.2 31.2 29.0 32.8
Siamdiff_atom 26.6 25.2 21.6 28.1 27.8 30.1
@@
Our Dataset ESM2_150M 19.5 19.6 16.6 20.5 18.9 22.0
GearNe 38.9 39.9 35.8 41.0 36.4 41.3
Siamdiff 37.0 37.7 31.0 40.7 35.3 40.3
Vabs-Net 58.3 60.1 52.0 58.8 56.3 64.1
The IoU results of prediction for the binding site of small molecules can be seen in the following table. We will add these results to our appendix. Results show that IoU drops slightly.
Table 8: Results of small molecules vary with the number of residues.
number of residues 0-200 200-399 400-599 600-799 800-999
Average IoU 0.663 0.675 0.612 0.501 0.614
Appendix B Vector Encoding.
In most protein pretraining models, distance is frequently utilized to encode protein structure. Theoretically, given KNN=30 (as detailed in our paper) as input, we should be capable of reconstructing a protein chain’s configuration (referring to the protein’s shape or the 3D coordinates of its atoms within a specific coordinate system.). Multidimensional scaling represents a widely employed algorithm for translating distances between each pair of n objects in a set into a configuration of n points, which are then mapped into an abstract Cartesian space. The computational complexity of this process is $O(N^{3})$, where N represents the number of points. Consequently, as the number of atom nodes increases, reconstructing the configuration of a protein becomes more challenging. Notably, the number of atoms is typically eight times greater (based on our pretraining dataset) than the number of residues, which suggests that distance may not be sufficient to encode all-atom configuration. This motivates us to add edge direction features to edges.
Appendix C More Detailed Information About Model Architecture.
@@
Vabs-Net 64.1 0.900
No SMPC 63.0 0.876
No SMPC SO(3) 60.9 0.874
@@
Throughout the fine-tuning process, with the exception of molecular docking, all parameters remained unfrozen. Additionally, in conventional practice, two evolutionary conversation profiles (PSSM and HMM) are traditionally utilized in DNA and RNA binding site prediction (Xia et al., 2021). A more comprehensive description of the model is provided in Table 11. The sizes of DNA and RNA molecules are considerably larger than those of small molecules, necessitating the selection of a k-nearest neighbor (knn) value of 90 to encompass long-range information. In contrast, the pre-training model maintains a knn value of 30. Due to the restricted size of the training set for DNA and RNA binding site prediction, we utilized a model with 256-node dimensions.
@@
Appendix E More Detailed Information About Equibind Training.
@@
Appendix F Computational Efficiency
All training experiments are conducted with 16 Tesla A100. For the downstream task, we choose small molecule binding site prediction. All inference experiments are conducted with 8 V100 on COACH420 test set(409 samples).
Table 11: Time Spent to Pre-train, Finetune, Inference Our All-atom Version and Residue-level Version vabs-net.
Settings Atom PT Residue PT Atom FT Residue FT Atom Infer Residue Infer
Time 4d16h 2d2h 8h 3.5h 22s 7s
Epochs 190 190 50 50 1 1